what is the probability that a market inspector will discover at most three violations of the public health code


1
Expert's answer
2021-05-21T11:09:26-0400

The probability that a market inspector will discover at most three violations of the public health code "= P(X \le 3) = P(X=0)+P(X=1)+P(X=2)+P(X=3)"